OpenStudy (anonymous):

Write in equation in standard-form of an equation for the line that passes through (7,-3) and has a y-intercept of 2

OpenStudy (shamil98):

y = mx + b -3 = 7m + 2 solve for m first.

OpenStudy (anonymous):

M=-5/7

OpenStudy (anonymous):

@shamil98 help